find the equation of cone generated by lines drawn through the origin to meet the circle through the points (3,0,0) ,(0,2,0),(0,0,1)
1
Expert's answer
2020-09-21T14:37:44-0400

Let find a plane S containing points A(3,0,0) ,B(0,2,0),C(0,0,1):

"\begin{vmatrix}
 x-x_a & y-y_a & z-z_a\\
 x_b-x_a & y_b-y_a & z_b-z_a\\ 
x_c-x_a & y_c-y_a & z_c-z_a\\
\end{vmatrix}" "= 0"


"\begin{vmatrix}

 x-3 & y-0 & z-0\\
 0-0 & 2-0 & 0-0\\ 
0-0 & 0-0 & 1-0\\

\end{vmatrix}" "=0"


Plane S equation:

"2*x+3*y+6*z-6=0"


Let find a height OH of a cone. OH is perpendicular to the plane S and passes through point O(0,0,0)

"\dfrac{x-0}{2}=\dfrac{y-0}{3}=\dfrac{z-0}{6}"

Let us find the equation of the straight line passing through the origin and point A

As a result, the parametric equation of the straight line D was obtained

"x=3*t"

Find the angle between the straight lines A and D

"\phi = 73.381°"

Length of line OH is 0.85:

"tg(\phi)=\dfrac{HA}{0.85}"

"HA = 2.78"

Equation of cone

"\dfrac{x^2}{2.78^2}+\dfrac{y^2}{2.78^2}-\dfrac{z^2}{0.85^2}=0"